40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jan 28 18:03:53 2021, shruti, DailyProgress, PSOMA, cavity locking Path.pdf
    Reply  Mon Feb 1 18:25:47 2021, shruti, DailyProgress, PSOMA, cavity locking 
       Reply  Wed Feb 3 10:56:54 2021, shruti, Update, PSOMA, cavity locking 
          Reply  Thu Feb 4 16:19:35 2021, shruti, Update, PSOMA, cavity alignment 5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg
             Reply  Thu Feb 4 22:27:43 2021, rana, Update, PSOMA, cavity alignment IMC_HOManalysis.ipynb
                Reply  Tue Feb 9 11:14:48 2021, shruti, Update, PSOMA, cavity alignment 
                   Reply  Tue Feb 16 11:30:14 2021, shruti, Update, PSOMA, Beat recovered, fiber organization, alignment F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg069D0D53-9111-4B33-96AD-05B907C4BBCB.jpegBeat_low_freq_oscillations.mp4
Message ID: 2632     Entry time: Tue Feb 9 11:14:48 2021     In reply to: 2631     Reply to this: 2646
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: cavity alignment 

I removed some of Zach's optics and a large post from the south west corner of the table and placed the Vitek monitor.

I also placed a HWP and PBS in the transmission path to split the light between the transmission PD and the Vitek camera. Hunting for a suitable DC power supply for the camera.

ELOG V3.1.3-